OpenStudy (anonymous):

6/25 as a decimal and also 0.605 as a fraction in simplest form

OpenStudy (anonymous):

6/25 = 24/100 = 0.24 0.605 = 605/1000 = 121/200

OpenStudy (anonymous):

you ok with that tron?

OpenStudy (anonymous):

yep XD